• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 12:23:49 on December 12,1999

    Entry number 30853


    keyword=trigger problems, solutions

    I got called in because there were no T1 or T3 in datastream,
    only T5. The reason was that the "AND" module which forms
    T5 was set in "OR" mode. This was because the CAMAC crate
    had been turned off (Thursday, I think), and the download software
    didn't communicate with this module. Going to the old software
    version fixed this. This will be the default for now. It seems fairly
    certain that other modules (MLU and delays) did indeed get downloaded.
    So the implication is that the data are ok, but one has to decide if
    T5 was actually T1 or T3 by looking at population of detectors.
    Below is a raw TC spectra for E-arm and H-arm TDC in run 1225,
    one of the runs affected. It looks ok. Peak at 225 on E-arm is for
    approx 600 MeV/C protons, and the false peak at 125 is normal; it
    is due to triggers that have no E-arm retimnig, hence retiming is
    "forced" to ensure DAQ synch. But the H-arm spectra happens to
    be immune from this oddity. The only negative implication I can
    think of is if one wants to work out deadtime corrections separately
    for each trigger type -- that would be impossible. Anyway, fixed now.

    FIGURE 1